OpenStudy (anonymous):

According to Josh’s measurement, the distance from his house to the baseball practice field is 4 km. The actual distance is 3.68 km. What was the percent error in Josh’s measurement? Round to the nearest tenth of a percent, if needed. A. 0.3% B. 8% C. 8.7% D. 32%

OpenStudy (legends):

When im doing the math its telling me 0.92

OpenStudy (legends):

Oh ok was doing it wrong

OpenStudy (anonymous):

What did you get for your answer?

OpenStudy (anonymous):

Take 4 and subtract 3.68 .

OpenStudy (anonymous):

Ok this problem works like this. I am going to use counting numbers, which makes it simpler I think. Lets say we measured 4 but the actual measurement was 2 . First fine the difference 4 - 2 = 2 Now what percent of 4 is 2. This is like saying what fraction of 4 is 2 2 / 4 = 1/2 = .5 = 50% The problem you have works the same way 4 - 3.68 = .32 Can you finish the rest?